OpenStudy (anonymous):

The sides of a right triangle are consecutive even integers. Find the length of each side. Should i use the Pythagorean theorem here?

OpenStudy (paxpolaris):

sure...

OpenStudy (paxpolaris):

let the smallest side be x

OpenStudy (anonymous):

how? x+2 + x+4 = x+6?

OpenStudy (anonymous):

x+2^2 + x+4 ^2 = x+6^2?

OpenStudy (paxpolaris):

your 3 sides will be x, x+2, x+4

OpenStudy (anonymous):

so x+4 will be my hypotenuse?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

Got it sir thanks :)